# Standardizing Generative Face Video Compression using Supplemental Enhancement Information

### v2

@@ III-C Encoding/Decoding Processes
The decoder receives the bitstream to reconstruct the face video. In particular, a VVC decoder with support for the GFV SEI message is utilized to decode the base picture and parse the GFV SEI message. Further, the parsed GFV SEI parameters and the VVC-decoded base pictures are jointly fed into the synthesis model (i.e., deep generative network) to render the reconstructed video. It should be mentioned that a software implementation that streamlines all encoding and decoding processes has been proposed and made available by JVET GFVC AhG for further experimentation [37]. Decoders that do not support the GFV SEI message would simply decode and display the result of the ordinary standardized core hybrid decoding process.
@@ IV-A Experimental Settings
IV-A1 Implementation Details
The GFV SEI message has been included in the working draft of VSEI version 4 [12] and its implementation on top of the VVC reference software has been made available in the VTM SEI branch 111VTM Software with GFV SEI Implementation. In addition, as the output of the JVET GFVC AhG activities, a GFVC software tool has been developed and made available on a GFVC git branch 222JVET AhG16 GFVC Software Tools. Additionally, python scripts are provided to link the GFVC software tool with the GFV SEI message implementation, such that the entire GFVC encoding/decoding processes using the GFV SEI approach can be executed in a streamlined manner.
@@ IV-A2 Testing Conditions
Experiments of GFVC using a GFV SEI message implementation were conducted under the JVET GFVC AhG test conditions [9]. Four representative GFVC algorithms (i.e., CFTE [4], FV2V [3], DAC [2] and FOMM [29]) and a lightweight GFVC version CFTE-Lite are selected as testing algorithms. All testing sequences 333JVET AhG16 Testing Sequences are available in two resolutions 256256 and 512512. Herein, these test sequences span considerable diversity of face video content in terms of age, gender, ethnicity, appearance, expression, head position, head motion, camera motion, background, etc. They are pre-processed (i.e., cropped) from open-source high-quality high-resolution video, and therefore have smaller resolutions relative to the original video and have squared aspect ratio. The detailed experimental settings are listed as follows:
For an anchor, VTM-22.2 reference software for VVC is used under low-delay B (LDB) configuration with quantization parameter (QP) set to 37, 42, 47 and 52.
For the tests, the first picture in each sequence is taken as the base picture and coded as an I picture using VTM-22.2 with QP values of 22, 32, 42 and 52. All other pictures are coded with the SEI-based GFVC approach, where the compact feature parameters are signaled via the GFV SEI message.
@@ IV-B Performance Comparisons
Fig. 2 shows that promising rate-distortion performance (i.e., Rate-DISTS and Rate-LPIPS) can be achieved with the proposed SEI-based GFVC approach compared to the traditional VVC standard in higher bitrate ranges for the 256256 resolution. It can be also seen that with the increase of face video resolution to 512512, the proposed SEI-based GFVC approach can achieve coding performance benefits for a wider range of bit rates. This is because the overhead of sending facial parameters via SEI messages becomes smaller relative to the VVC-coded base pictures as the video resolution increases. Such performance results confirm that SEI-based GFVC can provide promising coding performance for face video with resolutions at or above standard definition television resolution.
Fig. 3 provides subjective quality comparisons using VVC and the proposed SEI-based GFVC approaches at similar bit-rates. At ultra-low bitrates, VVC reconstructed face images are seriously distorted, whereas the proposed SEI-based approaches can still reconstruct vivid faces with higher fidelity and provide much better texture details.
@@ IV-C Functionalities Examples
In addition to ultra-low bandwidth video communication, this SEI-based GFVC approach enables various purposes like user-specified animation/filtering and metaverse-related applications, with capabilities that cannot be directly supported by the core VVC codec. As illustrated in Fig. 4, the proposed approach gives the sender control over the quality of animated or filtered face pictures, and it can also support parameter modification in the GFV SEI message for expression or posture interactions in a virtual world.
IV-D Complexity Analysis
IV-E Latency Analysis
Table III provides the average bit ratio that is used for the base picture and all pictures at different base picture QP settings. It can be revealed that the bit consumption of SEI-based GFVC approaches is more susceptible to the QP and resolution of base picture instead of GFV SEI messages. Specifically, as the QP of the base picture decreases or the resolution of the base picture increases, the bit ratio of base picture v.s. all pictures will both gradually increase.
In addition, Fig. 5 shows picture-level bit fluctuation for SEI-based GFVC approaches for two test sequences with 250 frames but different resolutions. It should be noted that the bit consumption of base picture is not counted but the bit consumption of its SEI messages is counted. As illustrated in Fig. 5, the proposed SEI-based GFVC approaches can stably allocate coding bits to all pictures without any obvious bit fluctuation. As such, this proposed SEI-based GFVC approach can well support ultra-low bitrate communication.
| SEI-based GFVC Methods | Resolution | GFVC Bit Ratio (Base Picture v.s. All Pictures) | |||||||||
@@ V Conclusions and Discussions
Generative AI has dramatically changed the acquisition of digital content and driven forward the innovative progress of visual data compression. However, these promising generative coding technologies face the challenge of providing a standardized way to represent various feature formats, thus greatly limiting their practical deployment and breadth of applications. To address this, this paper discusses how to define a unified high-level syntax structure for different GFVC representations and insert it into a VVC bitstream as SEI messages, and can incorporate different GFVC algorithms in a standardized workflow. The simulation results demonstrate the effectiveness of the proposed SEI-based GFVC approach in delivering promising performance and diverse functionalities.
As for potential future work related to GFVC technique and its standardization, more robust model and stable reconstruction quality may improve the feasibility of practical applications. Besides, it is necessary to further reduce model complexity especially for the GFVC decoder and reduce deployment cost especially on mobile devices. Another promising direction is to systematically investigate the quality evaluation of generative coding, since the existing pixel-level or perceptual-level quality measures are not entirely suitable to reflect the quality of reconstructed video. Furthermore, it is envisioned that the proposed approach could inspire the development of future generative video coding technologies beyond face video. In particular, it is straightforward to extend the proposed standardization scheme from talking face video to human moving body video or even some natural-scene video that possess strong prior knowledge in structure of motion and texture, which can facilitate the generative video coding towards diverse scenarios and collaborated communication.

### v3

@@ III-D Encoding/Decoding Processes
The decoder receives the bitstream to reconstruct the face video. In particular, a VVC decoder with support for the GFV SEI message is utilized to decode the base pictures and parse the GFV SEI message. Since the GFVC synthesis is expected to proceed in 4:4:4 RGB format, the decoded base pictures should execute a color format conversion prior to the synthesis Further, the parsed GFV SEI parameters and the converted base pictures are jointly fed into the synthesis model (i.e., deep generative network) to render the reconstructed video. It should be mentioned that a software implementation that streamlines all encoding and decoding processes has been proposed and made available by JVET GFVC AhG for further experimentation [60]. Decoders that do not support the GFV SEI message would simply decode and display the result of the ordinary standardized core hybrid decoding process.
@@ IV-A Experimental Settings
IV-A1 Implementation Specifics
The GFV SEI message has been included in the working draft of VSEI version 4 [18] and its implementation on top of the VVC reference software has been made available in the VTM SEI branch 111VTM Software with GFV SEI Implementation. In addition, as the output of the JVET GFVC AhG activities, a GFVC software tool has been developed and made available on a GFVC git branch 222JVET AhG16 GFVC Software Tools. Additionally, Python scripts are provided to link the GFVC software tool with the GFV SEI message implementation, such that the entire GFVC encoding/decoding processes using the GFV SEI approach can be executed in a streamlined manner. As for model training, each GenerativeNN (i.e., GFVC synthesis model) is trained on mixed data from VoxCeleb [61] and CelebV [62] training dataset for 100 epochs, whilst each TranslatorNN is trained on face features of VoxCeleb [61] training sets for 400 epochs. It should be noted that these neural models and network interfaces have been adopted in the JVET GFVC AhG activities and available in the JVET GFVC software tools.
@@ IV-A2 Testing Conditions
Experiments of GFVC using a GFV SEI message implementation were compared to the VVC codec according to the GFVC AhG test conditions [15], the adjusted VVC anchor and the latest learning-based video codec DCVC-FM [63]. Four representative GFVC algorithms (i.e., CFTE [6], FV2V [5], DAC [2] and FOMM [51]) and a lightweight GFVC version CFTE-Lite are selected as testing algorithms. All testing sequences 333JVET AhG16 Testing Sequences are available in two resolutions 256256 (Class A&B) and 512512 (Class C&D), where the total frame number of ClassA&C and ClassB&D sequences is 250 and 125, respectively. As shown in Fig. 4, these test sequences span considerable diversity of face video content in terms of age, gender, ethnicity, appearance, expression, head position, head motion, camera motion, background, etc. They are pre-processed (i.e., cropped) from open-source high-quality high-resolution video, and therefore have smaller resolutions relative to the original video and have squared aspect ratio. The detailed experimental settings are listed as follows:
For the VVC anchor, VTM-22.2 reference software for VVC is used under low-delay B (LDB) configuration with quantization parameter (QP) set to 37, 42, 47 and 52.
For DCVC-FM, low delay configuration is used with quantization factor set to 30, 20, 10 and 0.
For the SEI-based GFVC tests, the first picture in each sequence is taken as the base picture and coded as an I picture using VTM-22.2 with QP values of 22, 32, 42 and 52. All other pictures are coded with the SEI-based GFVC approach, where the compact feature parameters are signaled via the GFV SEI message. It should be noted that in our experimentation, the coding frequency of the base picture is 1 time, i.e., the base picture is not additionally updated during the whole encoding process.
To obtain more aligned bitrate ranges and reference quality with the GFVC algorithms, the QP settings for the VVC anchor (LDB mode) have been further adjusted. By doing this, the bitrates of the adjusted VVC anchor are effectively reduced, creating a larger range overlap with the proposed SEI-based GFVC method. The corresponding QP settings for the two testing resolutions are listed: 1) For the 256256 resolution, the QP of an I picture is set at and the QP of the subsequent frames are correspondingly fixed at . 2) For the 512512 resolution, the QP of an I picture is set at and the QP of the subsequent frames are correspondingly fixed at .
@@ IV-B Performance Comparisons
As illustrated in Fig. 5, we provide the rate-distortion performance comparisons among VVC, DCVC-FM and the proposed SEI-based GFVC apporach. Experimental results show that the proposed SEI-based GFVC approach can achieve promising rate-distortion performance in perceptual quality metrics like DISTS/LPIPS but perform poorly in pixel-level quality metrics like PSNR compared to the VVC codec according to the GFVC AhG test conditions, the adjusted VVC anchor and DCVC-FM. This result is consistent with the fact that the generative compression is designed in the feature domain rather than the pixel-level domain, so the reconstruction quality cannot achieve pixel-level alignment [66, 67].
In terms of the rate-distortion performance of the original VVC anchor and the adjuested VVC anchor, it can be observed that at the resolution of 256256, the adjusted VVC anchor shows a small performance improvement. However, as the resolution increases to 512512, the performance advantage in the high bitrate range completely disappears, and the adjusted VVC anchor underperforms the original VVC anchor. As such, it can be concluded that the GFVC testing conditions from JVET-AJ2035 [15] are relatively fair, and do not intentionally degrade the performance of the VVC codec to highlight the performance advantages of GFVC. Compared to both anchors and the DCVC-FM algorithm, our proposed SEI-based GFVC approach can achieve promising compression performance in higher bitrate ranges for the 256256 resolution. When the face video resolution is increased to 512512, the proposed SEI-based GFVC approach can achieve coding performance benefits for a wider range of bit rates. This is because the overhead of sending facial parameters via SEI messages becomes smaller relative to the VVC-coded base pictures as the video resolution increases. Such performance results confirm that SEI-based GFVC can provide promising coding performance for face video with resolutions at or above standard definition television resolution. Moreover, Fig. 5 shows the compression efficiency of these SEI-based GFVC methods. In particular, DAC achieves superior compression efficiency compared to FV2V and CFTE, with FOMM demonstrating the least effective performance. This hierarchical ranking directly demonstrates that DAC produces the most compact 2D keypoints representation to achieve the best compression efficiency among these tested GFVC methods.
Fig. 6 provides subjective quality comparisons using VVC, DCVC-FM and the proposed SEI-based GFVC approaches at similar bit-rates. At ultra-low bitrates, VVC-reconstructed or DCVC-FM-reconstructed face images suffer from severe artifacts like blockiness, blurring, and loss of fine facial features due to aggressive compression. In contrast, the proposed SEI-based approaches can still reconstruct vivid faces with higher fidelity and provide much better texture details. This stark disparity underscores the efficacy of the proposed SEI-based GFVC approaches in mitigating quality degradation for face video at extremely low bitrate ranges. Moreover, it can be observed that SEI-based GFVC methods sometimes result in imprecise reconstruction of local details like mouth and eye movements. For example, FV2V exhibits deviations in reconstructing the mouth region, while CFTE shows inaccuracies in the eye area, both of which are less effective than DAC in achieving accurate local reconstruction.
@@ IV-C Functionalities Examples
In addition to ultra-low bandwidth video communication, this SEI-based GFVC approach enables various purposes, such as user-specified animation/filtering and metaverse-related applications, with capabilities that cannot be directly supported by the core VVC codec. As illustrated in Fig. 7, the proposed approach gives the sender control over the quality of animated or filtered face pictures, and it can also support parameter modification in the GFV SEI message for expression or posture interactions in a virtual world.
In particular, using a virtual anime face for texture reference, a series of GFV SEI messages indicating realistic facial motion and expression can animate this reference image to generate face frames. This technology offers virtual character-based video animation and filtering, safeguarding user privacy in contexts like virtual live-streaming and secure video chatting. Moreover, the facial parameters embedded within the GFV SEI message correspond distinctly to various facial features and motions, such as mouth movements, eye blinks, head rotations, translations, and locations. This facilitates semantic-level controls and friendly interactions over the reconstructed face video, opening up exciting prospects for future video conferencing and metaverse-related activities.
Resolution | GFVC Bit Ratio (Base Picture v.s. All Pictures) | |||||||||
@@
IV-D Complexity Analysis
IV-E TranslatorNN Interoperability Analysis
To evaluate the effectiveness of TranslatorNN, we measure the RD performance of translated GFVC systems, where the encoder and decoder are equipped with different GFVC models utilizing different feature representations. The detail discussion and analysis of TranslatorNN can be referred to [46, 57, 68]. Herein, we take the CFTE algorithm as a test model to validate the feasibility of TranslatorNN in ‘CFTEtoAny’ scenario. Specifically, CFTE [6] is used as the encoder and DAC [2]/FOMM [51]/FV2V [5] are used as decoder, respectively. As illustrated Fig. 8, these translated GFVC systems like “CFTEtoDAC”, “CFTEtoFOMM” and “CFTEtoFV2V” show promising rate-DISTS performance and outperform VVC in a large margin. In other words, TranslatorNN can enable interoperability between different types of GFVC algorithms and avoid mismatch between the GFVC encoder and decoder. Moreover, TranslatorNN provides possibility to further improve RD performance. For example, for FOMM decoding system, the RD performances can be largely improved when a CFTE encoding system is utilized for less bit-rate consumption. The subjective quality of translated GFVC results and their original results are shown in Fig. 9. It can be observed that, with the aid of TranslatorNN, GFVC decoder can achieve vivid generations despite GFVC encoder mismatches.
IV-F Base-picture Coding Frequency Analysis
In the GFVC framework, the decision to update the base picture can be modeled as an optimization problem. Maintaining a fixed base picture achieves minimal bitrate consumption, but may degrade texture reference and reconstruction quality over time due to temporal variations. Conversely, periodic base picture updates incur higher coding costs but improve reconstruction fidelity.
To better discuss this point, we have added additional experimental analysis regarding base-picture coding frequency. Specifically, we use the CFTE algorithm as a test model and increase the base-picture encoding frequency from one to two and three times, and finally compare their RD results. As shown in Fig. 10, when the base-picture encoding frequency increases, the reconstruction of subsequent inter frames could be provided with better texture references, which can improve the objective quality of the reconstructed frames. However, the bit cost of increasing the base picture encoding frequency is huge, and it is difficult to achieve a good rate-distortion trade-off. Moreover, as reported in [2, 69, 54], compared with directly dividing the video into different GOPs to obtain the base picture, dynamically deciding which video frame is the base picture based on the temporal motion changes may be more intelligent and can facilitate the RD performance optimization. Thus, there is room for improvement in exploring the base picture update mechanism within the GFVC framework as a direction for future work.
IV-G Latency Analysis
Table III provides the average bit ratio that is used for the base picture and all pictures at different base picture QP settings. It can be revealed that the bit consumption of SEI-based GFVC approaches is more susceptible to the QP and resolution of base picture instead of GFV SEI messages. Specifically, as the QP of the base picture decreases or the resolution of the base picture increases, the bit ratio of base picture v.s. all pictures will both gradually increase. In addition, Fig. 11 shows picture-level bit fluctuation for SEI-based GFVC approaches for two test sequences with 250 frames but different resolutions. It should be noted that the bit consumption of base picture is not counted but the bit consumption of its SEI messages is counted. As illustrated in Fig. 11, the proposed SEI-based GFVC approaches can stably allocate coding bits to all pictures without any obvious bit fluctuation.
It should be mentioned that the proposed SEI-based GFVC approach follows a traditional low-delay setting in which the coding order and the display order are the same, and thus can well ensure high efficiency and low latency in video communication because SEI usually refers to additional data or features aimed at enriching the primary content, and these pieces of information are commonly transmitted and processed with high efficiency and low latency. As for the possible latency introduced by base-picture bitrate fluctuation, well-known techniques such as efficient buffer management, adaptive bitrate algorithms and network optimization can be employed. Such practices have long been successfully deployed in video conferencing applications to reduce the impact of bitrate fluctuations on latency in real-time video applications and are equally applicable to GFVC.
@@ V Conclusions and Discussions
Generative AI has dramatically changed the acquisition of digital content and driven forward the innovative progress of visual data compression. However, these promising generative coding technologies face the challenge of providing a standardized way to represent various feature formats, thus greatly limiting their practical deployment and breadth of applications. To address this, this paper discusses how to define a unified high-level syntax structure for different GFVC representations and insert it into a VVC bitstream as SEI messages, and can incorporate different GFVC algorithms in a standardized workflow. The simulation results demonstrate the effectiveness of the proposed SEI-based GFVC approach in delivering promising performance and diverse functionalities. The proposed GFV and GFVE SEI messages will be included in a new version of the VSEI standard that is expected to be finalized in the near future, and new versions of major video coding standards including VVC, HEVC and AVC will provide interface to enable the inclusion of GFV/GFVE SEI messages into VVC-, HEVC-, and AVC-coded bitstreams. This reflects the successful collaboration effort of multiple companies and research institutes to standardize generative video coding techniques within the JVET.
As for potential future work related to GFVC technique and its standardization, more robust model and stable reconstruction quality may improve the feasibility of practical applications. Besides, it is necessary to further reduce model complexity especially for the GFVC decoder and reduce deployment cost especially on mobile devices. The possible latency/delay problems in the real-time system also need to be solved to ensure the smoothness and stability of the playback. Another promising direction is to systematically investigate the quality evaluation of generative coding, since the existing pixel-level or perceptual-level quality measures are not entirely suitable to reflect the quality of reconstructed video. Furthermore, it is envisioned that the proposed approach could inspire the development of future generative video coding technologies beyond face video. In particular, it is straightforward to extend the proposed standardization scheme from talking face video to human moving body video or even some natural-scene video that possess strong prior knowledge in structure of motion and texture, which can facilitate the generative video coding towards diverse scenarios and collaborated communication.